# Modeling nonlinear scales with COLA: preparing for LSST-Y1

### v1

@@ III Results for LSST-Y1 Simulated Analysis
Refer to captionRefer to caption
Figure 4: Confidence contours (68% and 95%) of the parameters $\Omega_{m}$, $S_{8}$, and $w$ when the nonlinear matter power spectrum is computed using either ee2 (pink filled contours) or the baseline $w$CDM COLA emulator (dashed blue lines), all using the most aggressive Cutoff 3. In the top panel we place the fiducial cosmology at the reference cosmology of the emulator. This yields a low $\rm{FoB}~{}=0.11$, indicating that only the baseline emulator is needed when the reference is optimally positioned inside the posterior. However, the bottom panel shows that this same emulator produces large biases in regions far from the reference, yielding an $\rm{FoB}~{}=5.34$. These biases correlate strongly with $\sigma_{8}$, with the fiducial value used in the bottom panel ($\sigma_{8}=0.93$) being significantly greater than that of the ee2 reference cosmology ($\sigma_{8}=0.82$).
@@ III.1 Default-Precision Settings Using $N_{\mathrm{refs}}=1$
Nevertheless, given that in general the high-likelihood region is not known a priori, it is desirable to have an emulator that obtains accurate constraints throughout the whole parameter space. Hence, we perform MCMCs using fiducial cosmologies in which two cosmological parameters at a time are significantly shifted from the ee2 reference values, in order to reassess the agreement between the baseline COLA emulators and ee2 in these regions of the prior. We begin these tests by considering all 4 combinations of high and low $\Omega_{m}$ and $A_{s}$ values, and all 4 combinations of high and low $\Omega_{m}$ and $n_{s}$ values, according to Table 5. We then repeat the previous analysis using the baseline COLA emulator with Cutoff 3 with these 8 fiducial cosmologies. The 2D contours corresponding to the fiducial cosmology $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$, the most biased of these 8 cases, are shown in the bottom panel of Fig.  4 where we now see a strong disagreement between ee2 and COLA constraints. Specifically, the full $\rm{FoB}=5.34$ when the fiducial was shifted to $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$. However, in a test where the emulator’s single reference was set to the fiducial values of $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$, the FoB was restored to a low value of $0.40$, near the target of 0.3.
More comprehensively, the 1D bias metric of Eq. 27 between the baseline COLA emulator and ee2 for all 8 of the aforementioned fiducial cosmologies, and all three angular scale cuts, is shown in Fig. 5 for the parameters $\Omega_{m}$, $S_{8}$ and $w$. The fiducial cosmologies are listed in decreasing order of $\sigma_{8}$ as it displays a clear correlation with all parameter biases, especially for fiducial $\sigma_{8}$ values larger than the reference value $\sigma_{8,\rm ref}=0.82$. The largest 1D biases in $S_{8}$ between COLA and ee2 are for the fiducial cosmology $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$, which has $\sigma_{8}=0.93$. We also notice that $S_{8}$ tends to be the most biased parameter, as it is the most constrained by cosmic shear surveys. For the $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$ cosmology, the baseline emulator produces 1D biases in $S_{8}$ of $0.32\sigma_{S_{8}}$, $0.82\sigma_{S_{8}}$ and $1.33\sigma_{S_{8}}$ using Cutoffs 1, 2, and 3, respectively. The 6D FoB in these cases is $1.12$, $3.18$ and $5.34$. However, the maximum FoB for Cutoff 1 occurred for the $(\Omega_{m}^{\uparrow},n_{s}^{\uparrow})$ cosmology, which has the second largest $\sigma_{8}=0.90$, producing an FoB of 1.30.
Refer to captionRefer to caption
@@ III.4 Enhanced-Precision Settings Using $N_{\rm refs}^{\Lambda\rm CDM}=101$
Table 7 shows the fraction of points with $|\Delta\chi^{2}|<1$ for both $N_{\rm refs}^{\Lambda\rm CDM}=101$ emulators for the two most biased fiducial cosmologies. Across both fiducial cosmologies and three scale cuts we see very comparable goodness-of-fit. We conclude that, even with additional references, the enhanced-precision settings can only slightly help mitigating the COLA emulator biases. Given the improvement of using $N_{\rm refs}^{\Lambda\rm CDM}=101$, in Appendix C we test whether marginalizing over differences in the COLA and ee2 data vectors can reduce the remaining biases.
@@ IV Conclusion
In this work, we have tested using COLA-based emulators as a computationally affordable alternative for modelling the nonlinear power spectrum. We achieved this by comparing their performance to the high-resolution EuclidEmulator2 in a $w$CDM cosmic shear analysis of LSST-Y1 simulated data, employing three nonlinear scale cuts. In order to improve the performance of the COLA emulators, we tested two general techniques: incorporating high-resolution reference samples from ee2 and refining the resolution settings of the COLA simulations. We found that optimally placing a single reference prediction ($N_{\rm refs}=1$) is enough to calibrate COLA training simulations in a desired region of the prior for even the most aggressive scale cut, leading to low bias. However, this approach suffers from large bias in distant regions of the prior, especially with significantly different $\sigma_{8}$ values from that of the reference cosmology.
@@
It must be acknowledged that the $w$CDM model is only a minor extension to $\Lambda$CDM, introducing only one additional parameter. For more exotic models with several additional parameters, any approach relying upon $\Lambda$CDM references may degrade somewhat in performance. Nevertheless, we expect that these results hold for beyond-$\Lambda$CDM models that modify the expansion history of our Universe, such as dynamical dark energy theories, as well as theories that modify the linear order growth of structure in a scale-independent way [132].
Additionally, the COLA input itself can also be further improved beyond refining the force and mass resolutions. It is known that Particle-Mesh based N-body codes suffer from a loss of power at small scales due to their inability of resolving the internal structure of halos. To bypass this, different alternatives have been proposed in the literature, such as the Potential Gradient Descent methods [133] and neural network corrections to simulations [134]. Another possible avenue would be to use field-level emulation techniques such as the ones presented in [135, 136]. However, as our main proposal is to use COLA as a viable alternative to model nonlinear scales in beyond-$\Lambda$CDM models, an obvious setback in the implementation of field-level methods, is that they require us to run a decent amount of high-resolution N-body simulations of a given model to act as a true model. Therefore, while field level techniques are an extremely appealing alternative, they lack the needed predictability for investigations of alternatives to the Standard Model.
@@ A.3 Neural Network
In order to prevent overfitting, we perform a $L_{1}-L_{2}$ regularization, also known as Elastic Net regularization, of the weights and biases. This encourages the training process to keep weights and biases to a low absolute value, limiting the complexity of the model. We keep $10\%$ of the training cosmologies as a validation set for early stopping: a technique that prevents overfitting by ensuring that the validation loss does not grow during the training. The training process is halted after 2500 epochs, with an initial learning rate of $10^{-3}$ which, starting from epoch $1500$, halves every $200$ epochs. Our neural network is implemented using the keras [^13] https://keras.io library, and the loss function is minimized using the adam optimizer [141].
@@ A.6 Testing Emulator Equivalence
We now test whether the three emulator techniques, namely Gaussian Process regression, neural networks, and Polynomial Chaos Expansion, are equivalent at the level of parameter inference. For this test, we ran MCMCs using the LSST-Y1 simulated cosmic shear data. We utilized the three emulated methods, where each emulator was trained with default-precision simulations and $N_{\rm refs}=1$. We use the ee2 reference cosmology as the fiducial, and employ a scale cut even more aggressive than Cutoff 3, with $\theta_{\mathrm{min}}=2.7^{\prime}$ for $\xi_{+}$ and $\theta_{\mathrm{min}}=8.7^{\prime}$ for $\xi_{-}$. Fig. 12 shows the confidence contours and marginalized posteriors of several cosmological parameters obtained with the three emulators. We see strong agreement in the 1D distributions, and we conclude that any of the three emulation techniques can be used interchangeably.
@@
Figure 14: Following Fig. 1, the comparisons of the boost from COLA to ee2, for the training set and validation sets, using the $N_{\rm refs}^{\Lambda\rm CDM}=500$ and $N_{\rm refs}^{\Lambda\rm CDM}=\infty$ methodology. For $N_{\rm refs}^{\Lambda\rm CDM}=500$ references we compute the boost using $B(k,z)=\tilde{B}(k,z)$ as in Eq. 5, while for $N_{\rm refs}^{\Lambda\rm CDM}=\infty$ references we use $B(k,z)=\bar{B}(k,z)$ as in Eq. 29. For these comparisons, we evaluate the respective equations for the boost by computing $B_{\rm COLA}(k,z)$ directly from simulations for the $w$CDM predictions, while using the NPCE emulator for the $\Lambda$CDM-projected predictions as we lack the requisite $\Lambda$CDM simulations.
@@
Appendix C Mitigating Biases Using Principal Component Analysis

### v2

@@ III Results for LSST-Y1 Simulated Analysis
Figure 4: Confidence contours (68% and 95%) of the parameters $\Omega_{m}$, $S_{8}$, and $w$ when the nonlinear matter power spectrum is computed using either ee2 (pink filled contours) or the baseline $w$CDM COLA emulator (dashed blue lines), all using the most aggressive Cutoff 3. In the top panel we place the fiducial cosmology at the reference cosmology of the emulator. This yields a low $\rm{FoB}~{}=0.11$, indicating that only the baseline emulator is needed when the reference is optimally positioned inside the posterior. However, the bottom panel shows that this same emulator produces large biases in regions far from the reference, yielding an $\rm{FoB}~{}=5.37$. These biases correlate strongly with $\sigma_{8}$, with the fiducial value used in the bottom panel ($\sigma_{8}=0.93$, $S_{8}=1.02$) being significantly greater than that of the ee2 reference cosmology ($\sigma_{8}=0.82$, $S_{8}=0.84$).
@@ III.1 Default-Precision Settings Using $N_{\mathrm{refs}}=1$
Nevertheless, given that in general the high-likelihood region is not known a priori, it is desirable to have an emulator that obtains accurate constraints throughout the whole parameter space. Hence, we perform MCMCs using fiducial cosmologies in which two cosmological parameters at a time are significantly shifted from the ee2 reference values, in order to reassess the agreement between the baseline COLA emulators and ee2 in these regions of the prior. We begin these tests by considering all 4 combinations of high and low $\Omega_{m}$ and $A_{s}$ values, and all 4 combinations of high and low $\Omega_{m}$ and $n_{s}$ values, according to Table 5. We then repeat the previous analysis using the baseline COLA emulator with Cutoff 3 with these 8 fiducial cosmologies. The 2D contours corresponding to the fiducial cosmology $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$, the most biased of these 8 cases, are shown in the bottom panel of Fig.  4 where we now see a strong disagreement between ee2 and COLA constraints. Specifically, the full $\rm{FoB}=5.37$ when the fiducial was shifted to $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$. However, in a test where the emulator’s single reference was set to the fiducial values of $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$, the FoB was restored to a low value of $0.40$, near the target of 0.3.
More comprehensively, the 1D bias metric of Eq. 27 between the baseline COLA emulator and ee2 for all 8 of the aforementioned fiducial cosmologies, and all three angular scale cuts, is shown in Fig. 5 for the parameters $\Omega_{m}$, $S_{8}$ and $w$. The fiducial cosmologies are listed in decreasing order of $\sigma_{8}$ as it displays a clear correlation with all parameter biases, especially for fiducial $\sigma_{8}$ values larger than the reference value $\sigma_{8,\rm ref}=0.82$. The largest 1D biases in $S_{8}$ between COLA and ee2 are for the fiducial cosmology $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$, which has $\sigma_{8}=0.93$. We also notice that $S_{8}$ tends to be the most biased parameter, as it is the most constrained by cosmic shear surveys. For the $(\Omega_{m}^{\uparrow},A_{s}^{\uparrow})$ cosmology, the baseline emulator produces 1D biases in $S_{8}$ of $0.32\sigma_{S_{8}}$, $0.82\sigma_{S_{8}}$ and $1.35\sigma_{S_{8}}$ using Cutoffs 1, 2, and 3, respectively. The 6D FoB in these cases is $1.12$, $3.18$ and $5.37$. However, the maximum FoB for Cutoff 1 occurred for the $(\Omega_{m}^{\uparrow},n_{s}^{\uparrow})$ cosmology, which has the second largest $\sigma_{8}=0.90$, producing an FoB of 1.27.
@@ III.4 Enhanced-Precision Settings Using $N_{\rm refs}^{\Lambda\rm CDM}=101$
Table 8 shows the fraction of points with $|\Delta\chi^{2}|<1$ for both $N_{\rm refs}^{\Lambda\rm CDM}=101$ emulators for the two most biased fiducial cosmologies. Across both fiducial cosmologies and three scale cuts we see very comparable goodness-of-fit. We conclude that, even with additional references, the enhanced-precision settings can only slightly help mitigating the COLA emulator biases. Given the improvement of using $N_{\rm refs}^{\Lambda\rm CDM}=101$, in Appendix E we test whether marginalizing over differences in the COLA and ee2 data vectors can reduce the remaining biases.
@@ IV Conclusion
In this work, we have tested using COLA-based emulators as a computationally affordable alternative for modelling the nonlinear power spectrum. We achieved this by comparing their performance to the high-resolution EuclidEmulator2 in a $w$CDM cosmic shear analysis of LSST-Y1 simulated data, employing three nonlinear scale cuts. In order to improve the performance of the COLA emulators, we tested two general techniques: incorporating high-resolution reference samples from ee2 and refining the resolution settings of the COLA simulations. We emphasize that the choice to use reference predictions from ee2 was made solely because it was the benchmark by which we were measuring accuracy of our COLA emulators, and we encourage future exploration into using other prescriptions as references. We found that optimally placing a single reference prediction ($N_{\rm refs}=1$) is enough to calibrate COLA training simulations in a desired region of the prior for even the most aggressive scale cut, leading to low bias. However, this approach suffers from large bias in distant regions of the prior, especially with significantly different $\sigma_{8}$ values from that of the reference cosmology.
@@
It must be acknowledged that the $w$CDM model is only a minor extension to $\Lambda$CDM, introducing only one additional parameter. For more exotic models with several additional parameters, any approach relying upon $\Lambda$CDM references may degrade somewhat in performance. Nevertheless, we expect that these results hold for beyond-$\Lambda$CDM models that modify the expansion history of our Universe, such as dynamical dark energy theories, as well as theories that modify the linear order growth of structure in a scale-independent way [139, 140, 141].
Additionally, the COLA input itself can also be further improved beyond refining the force and mass resolutions. It is known that Particle-Mesh based N-body codes suffer from a loss of power at small scales due to their inability of resolving the internal structure of halos. To bypass this, different alternatives have been proposed in the literature, such as the Potential Gradient Descent methods [142] and neural network corrections to simulations [143]. Another possible avenue would be to use field-level emulation techniques such as the ones presented in [144, 145, 146, 147]. However, as our main proposal is to use COLA as a viable alternative to model nonlinear scales in beyond-$\Lambda$CDM models, an obvious setback in the implementation of field-level methods, is that they require us to run a decent amount of high-resolution N-body simulations of a given model to act as a true model. Therefore, while field level techniques are an extremely appealing alternative, they lack the needed predictability for investigations of alternatives to the Standard Model.
@@ A.3 Neural Network
In order to prevent overfitting, we perform a $L_{1}-L_{2}$ regularization, also known as Elastic Net regularization, of the weights and biases. This encourages the training process to keep weights and biases to a low absolute value, limiting the complexity of the model. We keep $10\%$ of the training cosmologies as a validation set for early stopping: a technique that prevents overfitting by ensuring that the validation loss does not grow during the training. The training process is halted after 2500 epochs, with an initial learning rate of $10^{-3}$ which, starting from epoch $1500$, halves every $200$ epochs. Our neural network is implemented using the keras [^15] https://keras.io library, and the loss function is minimized using the adam optimizer [152]. The training process for all redshifts takes 2 hours without GPU acceleration, and the full emulator takes around 250MB of disk space.
@@ A.6 Testing Emulator Equivalence
We now test whether the three emulator techniques, namely Gaussian Process regression, neural networks, and Polynomial Chaos Expansion, are equivalent at the level of parameter inference. For this test, we ran MCMCs using the LSST-Y1 simulated cosmic shear data. We utilized the three emulated methods, where each emulator was trained with default-precision simulations and $N_{\rm refs}=1$. We use the ee2 reference cosmology as the fiducial, and employ a scale cut even more aggressive than Cutoff 3, with $\theta_{\mathrm{min}}=2.7^{\prime}$ for $\xi_{+}$ and $\theta_{\mathrm{min}}=8.7^{\prime}$ for $\xi_{-}$. Fig. 12 shows the confidence contours and marginalized posteriors of several cosmological parameters obtained with the three emulators. We see strong agreement in the 1D distributions, and we conclude that any of the three emulation techniques can be used interchangeably. However, we note that there are differences in the training time of the various emulators. Considering all principal components and redshifts, and using only CPUs, the GP emulators took $\sim 10$ minutes to train on 1 core, the NN took $\sim 2$ hours to train on 1 core, the PCE emulators took $\sim 10$ hours to train on 16 cores, and the NPCE emulators took an extra $\sim 6$ hours to train compared to the PCE.
@@
Figure 14: Following Fig. 1, the comparisons of the boost from COLA to ee2, for the training set and validation sets, using the $N_{\rm refs}^{\Lambda\rm CDM}=500$ and $N_{\rm refs}^{\Lambda\rm CDM}=\infty$ methodology. For $N_{\rm refs}^{\Lambda\rm CDM}=500$ references we compute the boost using $B(k,z)=\tilde{B}(k,z)$ as in Eq. 5, while for $N_{\rm refs}^{\Lambda\rm CDM}=\infty$ references we use $B(k,z)=\bar{B}(k,z)$ as in Eq. 29. For these comparisons, we evaluate the respective equations for the boost by computing $B_{\rm COLA}(k,z)$ directly from simulations for the $w$CDM predictions, while using the NPCE emulator for the $\Lambda$CDM-projected predictions as we lack the requisite $\Lambda$CDM simulations. For comparisons at higher redshift see Appendix C.
@@
Appendix C Boost Comparisons to ee2 at Higher Redshift
In this section we include comparisons of the post-processed boosts from our COLA simulations to those computed by ee2, at higher redshifts that more significantly impact the LSST simulated analyses performed in the paper (see Tab. 6 for the mean redshift of each galaxy source bin used). These comparisons are presented in Fig. 15 where we show the 68th percentile error in the boost at each wavenumber and redshift. From the figure we see that in all cases the disagreement at higher redshifts is either roughly consistent with the error at $z=0$ or noticeably decreased.
Figure 15: In this figure we show the 68th percentile error between the post-processed boost from our COLA training simulations and those from ee2, at each wavenumber and redshift, using many of the different approaches tested in this paper. The error is computed as the absolute value in the percentage difference in the boost $100\times|B/B_{\textsc{ee2}}-1|$, where $B$ is replaced by $\bar{B}_{\rm COLA}$ or $\tilde{B}_{\rm COLA}$ via Eqs. 4, 5, or  29 in the respective cases. The labels identify the model within which the training simulations are spread, the precision settings of the simulations, and the number and model of the references used. While the top row compares results for $N_{\rm refs}=1$, the middle row compares the $N_{\rm refs}=101$ cases showing the lack of degradation when confining the references to $\Lambda$CDM, and the bottom row shows the improvement due to increasing $N_{\rm refs}^{\Lambda\rm CDM}$.
Appendix D Generating Fiducial Data with the bacco Emulator
In the analysis presented in III, ee2 was used both in generating the fiducial data, as well as in modelling the data vector at the sampled points in the chains. This is true not only for chains in which ee2 was used directly to model the nonlinear power spectrum, but also for COLA emulator chains since ee2 was used to compute the reference predictions. While our original analysis only compared the MCMC constraints between the COLA emulators and ee2, it remains possible in theory that generating the fiducial data with another prescription could cause the COLA emulator and ee2 constraints to diverge more. This is an important consideration as real data from LSST will not be generated with ee2.
We briefly explore whether our results depend strongly on the choice of nonlinear prescription used in generating the fiducial data, by repeating previous tests with fiducial data generated with the bacco emulator. As it was determined in Sec. III that the $N_{\rm refs}^{\Lambda\rm CDM}=500$ and $N_{\rm refs}^{\Lambda\rm CDM}=\infty$ approaches were the most accurate, we repeat only tests similar to those presented in Fig. 11. However, since bacco emulator only covers values of $\sigma_{8}<0.9$, we generated fiducial data vectors at $\Omega_{m}=0.33$ while maintaining the other cosmological parameters. We also used ee2 to generate fiducial data for $z>1.5$ as this is outside of the bacco emulator’s range.
Fig. 16 shows the 1D bias measures of Eq. 27 between ee2 and cola emulators. Once again, we confirm that COLA emulation under the $N_{\rm refs}^{\Lambda\rm CDM}=500$ and $N_{\rm refs}^{\Lambda\rm CDM}=\infty$ approaches produced low bias relative to ee2, now with a different prescription for the fiducial data. For Cutoffs 1, 2, and 3 the worst FoB values produced by the $N_{\rm refs}^{\Lambda\rm CDM}=500$ emulator were 0.09, 0.23, and 0.50 respectively. Similarly the highest FoB values produced using the $N_{\rm refs}^{\Lambda\rm CDM}=\infty$ approach were 0.10, 0.16, and 0.38 for the three scale cuts, either meeting or near the target value of 0.3.
Figure 16: 1D biases between cola emulators and ee2 with fiducial data vectors generated with bacco. These test the robustness of the $N_{\rm refs}^{\Lambda\rm CDM}=500$ and $N_{\rm refs}^{\Lambda\rm CDM}=\infty$ results presented in Fig. 11. Note that the fiducial $\Omega_{m}$ value differs from past tests in order to bring the cosmology inside the bacco parameter space.
Cosmo. Emul. DP, 500 ($\Lambda$CDM) DP, $\infty$ ($\Lambda$CDM)
$\Omega_{m}=0.33$, $A_{s}^{\uparrow}$, $w^{\downarrow}$, C1 $95.0\%$ $98.0\%$
$\Omega_{m}=0.33$, $A_{s}^{\uparrow}$, $w^{\downarrow}$, C2 $77.8\%$ $87.5\%$
$\Omega_{m}=0.33$, $A_{s}^{\uparrow}$, $w^{\downarrow}$, C3 $43.0\%$ $52.0\%$
$\Omega_{m}=0.33$, $n_{s}^{\uparrow}$, $w^{\downarrow}$, C1 $94.0\%$ $96.6\%$
$\Omega_{m}=0.33$, $n_{s}^{\uparrow}$, $w^{\downarrow}$, C2 $74.3\%$ $87.1\%$
$\Omega_{m}=0.33$, $n_{s}^{\uparrow}$, $w^{\downarrow}$, C3 $50.3\%$ $57.8\%$
Table 10: Percentages of MCMC accepted points with the difference in $\chi^{2}$ between the COLA emulators and ee2 falling in the desired range of $|\Delta\chi^{2}|<1$, when the fiducial data vectors are generated with the bacco emulator.
Appendix E Mitigating Biases Using Principal Component Analysis